OpenStudy (anonymous):

I need some Math help please.... Indicate in standard form the equation of the line passing through the given points. R(3, 3), S(-6, -6)

OpenStudy (lgbasallote):

use the formula \[\Large y - y_1 = \frac{y_2 - y_1}{x_2 - x_1} (x-x_1)\] does that help?

OpenStudy (anonymous):

I got... Y-3=-9 over 9=X-3

OpenStudy (lgbasallote):

close... \[\Large \frac{y_2 - y_1}{x_2 - x_1} \implies \frac{-6 - 3}{-6 - 3} \implies \frac{-9}{-9} \implies 1\] so the equation should be something like y - 3 = 1(x-3) do you get it?

OpenStudy (anonymous):

Yeah. So that's the answer?

OpenStudy (lgbasallote):

not yet. simplify it

OpenStudy (anonymous):

Now I'm lost. I have no idea how to simplify it.

OpenStudy (lgbasallote):

what's 1(x-3)?

OpenStudy (anonymous):

I don't know. Distributive property?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

So.. is the answer 1x-3?

OpenStudy (lgbasallote):

right.. y - 3 = x- 3 now add 3 to both sides

OpenStudy (anonymous):

Alright man, thanks I appreciate the help!